OpenStudy (dls):

How far away from the surface of the Earth does the acceleration due to gravity become 4% of its value on the surface of the Earth?

OpenStudy (dls):

@experimentX !

OpenStudy (dls):

\[\LARGE \frac{4}{100}g=\frac{g}{(1+\frac{h}{R})^2}\] is this the equation?

OpenStudy (dls):

ah sorry for disturbing :P its done